# Understanding the Training and Generalization of Pretrained Transformer for Sequential Decision Making

### v1

@@ 4 Learned Decision Function as an Algorithm
There are simple remedies to the linear regret behavior of the optimal decision function, for example, using an $\epsilon$-greedy version of the transformer. Numerically, we do not observe such behavior for any experiment even without any further modification of the pretrained transformer. Proposition 4.4 provides a simple explanation for why one does not have to be too pessimistic. Intuitively, the pretrained model $\texttt{TF}_{\theta}$ can be viewed as the optimal decision function plus some random noise, and the random noise actually plays the role of exploration and helps the posterior to concentration. Proposition 4.4 involves two conditions. The first condition is that $\texttt{TF}_{\hat{\theta}}$ behaves closely to $\texttt{Alg}^{*}$ and the deviation between these two is captured by $\epsilon$. This condition can be justified by Figure 2. The second condition characterizes the KL divergence between the conditional distribution of the observation $o_{t}$ for any two environments. Note that this KL divergence critically depends on $\texttt{TF}_{\theta}$ because the ability to distinguish between two environments depends on the intensity of exploration of the underlying algorithm. These two conditions work more like a stylized model to analyze the behavior of the transformer. It is a largely simplified model; say, the parameter $\Delta$ should have a dependency on $t$ for many online learning algorithms such as UCB or Thompson sampling. Nevertheless, the two conditions give us a perspective to understand how the transformer works as an online algorithm. In particular, the transformer deviates from $\texttt{Alg}^{*}$ by an amount of $\epsilon$ and this deviation accumulates over time into the first term in the regret bound. On the other hand, this deviation gives an exploration, and the second term captures the regret caused during the concentration of the posterior onto the true environment. In Appendix D.6.1, we provide some justifications for the two conditions, and also establish $\Delta=\Theta(\epsilon)$ for linear bandits and dynamic pricing. We also note that the finiteness of $\mathcal{P}_{\gamma}$ is not essential. The KL divergence condition can be modified for a continuous $\mathcal{P}_{\gamma}$ and combined with a covering argument to obtain the bound accordingly.
@@ 5 Numerical Experiments and Discussions
We make the following observations on the numerical experiment. The transformer achieves a better performance than the structured algorithms for each task. This result is impressive to us in that the transformer is not replicating/imitating the performance of these benchmark algorithms but it discovers a new and more effective algorithm. This in fact shows the effectiveness of using the optimal actions $a_{t}^{*}$ as the target variable for training the transformer. Moreover, we attribute the advantage of the transformer against the benchmark algorithms to two aspects: prior knowledge from pretraining data and more greedy exploitation. First, the transformer well utilizes the pretraining data and can be viewed as tailored for the pretraining distribution. Comparatively, all the benchmark algorithms are cold-start. Second, the benchmark algorithms are all designed in ways to cater for asymptotic optimality and thus may sacrifice short-term rewards. Comparatively, the transformer behaves more greedy; just like $\texttt{Alg}^{*}$ is a decision function that does not explicitly encourage exploration. And such greedy behavior brings better reward when $T$ is small such as $T\leq 100.$ More numerical experiments are deferred to Section C.
Conclusion and limitations. In this paper, we study the pretrained transformer for sequential decision-making problems. Such problems enable using the optimal action as the target variable for training the transformer. We investigate the training and generation aspect of the transformer and identify an OOD issue that is largely ignored by the existing literature. Moreover, we interpret how the transformer works as an online algorithm through the lens of the Bayes-optimal decision function. The numerical experiments are encouraging in that the trained transformer learns a decision function better than all the benchmark algorithms. However, we point out that all the numerical experiments in this paper are conducted under small dimensions with $\text{dim}(X_{t})\leq 10.$ Such settings provide a testbed for us to understand the working mechanism of the trained transformer; but to make it more practically useful, scaling up its ability for high-dimensional settings is necessary. Our conjecture is that this will entails a longer training procedure as well as a potential larger model architecture.
Appendix A More Related Works
@@
A remarkable characteristic of pretrained transformers is their ability to perform in-context learning (ICL). Once pretrained on a vast corpus, transformers can solve new tasks with just a few examples, without updating their parameters. ICL has captured the attention of the theoretical machine learning community, leading to considerable efforts into understanding ICL from different theoretical perspectives [Xie et al., 2021, Akyürek et al., 2022, Von Oswald et al., 2023, Lin et al., 2023]. A stream of literature that is related to our work explains the transformer’s behavior as Bayesian inference. See Xie et al. [2021], Zhang et al. [2023b] for the Bayesian argument in regression context, and Lee et al. [2024] for the Bayesian analysis in decision-making. Another stream of literature aims to develop generalization bound or other types of convergence results. See Jeon et al. [2024], Bai et al. [2024] for such results in the regression context and Lin et al. [2023] in the decision-making context.
Pretrained transformer for RL.
The earliest effort of applying transformers to reinforcement learning lies in the area of offline RL [Janner et al., 2021, Chen et al., 2021]. By autoregressively maximizing the likelihood of trajectories in the offline dataset, this paradigm essentially converts offline RL to a supervised learning problem. When evaluating the policy, the actions are sampled according to the likelihood of PT conditioned on the so called “return-to-go”, which is commonly chosen to be the cumulative reward of “good” trajectories in the dataset. By doing so, PT retrieves information from trajectories in training data with similar return-to-go values and performs imitation learning to high-reward offline trajectories. Based on this observation, several works further examine the validity of return-to-go conditioning [Emmons et al., 2021, Ajay et al., 2022, Brandfonbrener et al., 2022] and propose alternative methods that improve this approach or extend to other settings [Laskin et al., 2022, Badrinath et al., 2024, Zheng et al., 2022, Wu et al., 2024]. For example, one of the drawbacks of the return-to-go conditioning is that it often fails in trajectory-stitching when offline data comes from sub-optimal policies.
@@ Appendix B Problem Examples for the General Setup in Section 2
For the general problem setup in Section 2, the following proposition relates the pretraining and the regret through the lens of loss function.
@@
$\mathrm{Regret}(\texttt{TF}_{\theta};\gamma)\leq C\cdot L\left(\texttt{TF}_{\theta};\gamma\right)$
holds for any $\texttt{TF}_{\theta}$ and $\gamma$. We have
@@
Proposition B.2 gives a first validity of the supervised pretraining approach. The surrogate property is a property of the loss function and it is not affected by the distribution or the decision function $\texttt{TF}_{\theta}.$ Specifically, it upper bounds the regret with the action prediction loss. It is also natural in that a closer prediction of the optimal action gives a smaller regret. One implication is that the result guides the choice of the loss function for different underlying problems.
@@
To prove Proposition B.2, without loss of generality, we assume the following mild conditions. First, for all these problems, the noise random variable $\epsilon_{t}$ has mean 00 and is independent of the action $a_{t}$. Second, for the linear bandits problem, the $L_{\infty}$ norm of the unknown $w$ is bounded by $D$. Third, for the dynamic pricing problem and the newsvendor problem, we assume the demand function is differentiable with respect to the action coordinate, and the first and second-order derivatives are $D$-Lipschitz.
@@
Now we prove Proposition B.2 part by part.
(a). For the multi-arm bandits problem, let the optimal arm is the first one. For $1\leq i\leq k$, let $\mu_{i}$ be the mean of the reward distribution for arm $i$, and denote by $\Delta_{i}=\mu_{1}-\mu_{i}$ the action gap for arm $i$. Let $\Delta_{\max}=\max\left\{\Delta_{1},\cdots,\Delta_{k}\right\}$ be the maximum action gap.
Since $\texttt{TF}_{\theta}(H_{t})$ can be viewed as a distribution over $\left\{1,\cdots,k\right\}$, we now rewrite $P(\texttt{TF}_{\theta}(H_{t})=i)$ as $p^{t}_{i}$ for notation simplicity. We claim that $\text{Regret}(\texttt{TF}_{\theta};\gamma)\leq\Delta_{\max}\cdot L(\texttt{TF}_{\theta};\gamma)$, where $L(\cdot,\cdot)$ is the cross entropy loss. To see this, observe that condition on $H_{t}$,
$\mathbb{E}\left[r\left(X_{t},a_{t}^{*}\right)-r\left(X_{t},a_{t}\right)|H_{t}\right]$ $=\mu_{1}-\sum_{i=1}^{k}p^{t}_{i}\mu_{i}=\sum_{i=2}^{k}p^{t}_{i}\Delta_{i}\leq\Delta_{\max}(1-p^{t}_{1})\leq-\Delta_{\max}\cdot\log(p^{t}_{1}),$
then take the expectation of both sides and take the sum from $t=1$ to $t=T$ of both sides completes the proof.
(b). For the dynamic pricing problem, we will show that $\text{Regret}(\texttt{TF}_{\theta};\gamma)\leq D\cdot L(\texttt{TF}_{\theta};\gamma)$, where $L(\cdot,\cdot)$ is the squared $L_{2}$ loss. From Taylor’s expansion, we have for any $a_{t}\in\mathcal{A}$,
$r(X_{t},a_{t})-r(X_{t},a^{*}_{t})\geq-D||a_{t}-a_{t}^{*}||_{2}^{2},$
which implies that $\mathbb{E}\left[r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]\leq D||a_{t}-a_{t}^{*}||_{2}^{2}.$ Then taking the sum from $t=1$ to $t=T$ gives the results.
[…]
$\mathbb{E}\left[r\left(X_{t},a_{t}^{*}\right)-r\left(X_{t},a_{t}\right)\right]=\mathbb{E}\left[w^{T}a^{*}-w^{T}a_{t}\right]\leq\mathbb{E}\left[||w||_{\infty}||a^{*}-a_{t}||_{1}\right]\leq D\cdot\mathbb{E}\left[||a_{t}^{*}-f_{\theta}(H_{t})||_{1}\right].$
[…]
$r(X_{t},a_{t}^{*})-r(X_{t},a_{t})$ $\leq\max\left\{h,l\right\}|a_{t}^{*}+d(X_{t},a_{t}^{*})-a_{t}-d(X_{t},a_{t})|$
$=\max\left\{h,l\right\}\left(\mathbf{1}+\nabla_{a}d(X_{t},\xi)\right)^{T}(a_{t}^{*}-a_{t})$
$\leq(D+1)\max\left\{h,l\right\}||a_{t}^{*}-a_{t}||_{1}.$
Then taking the expectation and taking the sum from $t=1$ to $t=T$ of both sides gives the result. ∎

### v2

@@ 4 Learned Decision Function as an Algorithm
There are simple remedies to the linear regret behavior of the Bayes-optimal decision function, for example, using an $\epsilon$-greedy version of $\texttt{Alg}^{*}$ (and $\texttt{TF}_{\hat{\theta}}$). Numerically, we do not observe such behavior for any experiment even without any further modification of the pre-trained transformer $\texttt{TF}_{\hat{\theta}}$. Proposition 4.4 provides a simple explanation for why one does not have to be too pessimistic. Intuitively, the pre-trained model $\texttt{TF}_{\hat{\theta}}$ can be viewed as the optimal decision function $\texttt{Alg}^{*}$ plus some random noise, and the random noise actually plays the role of exploration and helps the posterior to concentration. Proposition 4.4 materializes such an intuition and it involves two conditions. The first condition is that $\texttt{TF}_{\hat{\theta}}$ behaves closely to $\texttt{Alg}^{*}$ and the (reward) deviation between these two is captured by $\Delta_{\text{Exploit}}$. This condition can be justified by Figure 3. The second condition characterizes the KL divergence between the conditional distribution of the observation $o_{t}$ for two environments: the true testing environment $\gamma$ and any other one $\gamma^{\prime}\in\Gamma$. Note that this KL divergence critically depends on $\texttt{TF}_{\hat{\theta}}$ because the ability to distinguish between two environments depends on the intensity of exploration of the underlying algorithm. These conditions work more like a stylized model to understand the behavior of the transformer. It is a largely simplified model; say, the parameter $\Delta_{\text{Explore}}$ should have a dependency on $t$ for many online learning algorithms such as UCB or Thompson sampling. Nevertheless, the conditions give us a perspective to understand how the transformer works as an online algorithm. In particular, $\texttt{TF}_{\hat{\theta}}$ deviates from $\texttt{Alg}^{*}$ by an amount of $\Delta_{\text{Exploit}}$ in terms of the reward, and this deviation accumulates over time into the first term in the regret bound. On the other hand, this deviation also gives an exploration whose intensity is measured by $\Delta_{\text{Explore}}$, and the second term captures the regret caused during the concentration of the posterior onto the true environment. In Appendix D.7, we provide some further justifications for these conditions. We also note that the finiteness of $\Gamma$ is not essential. The KL divergence condition can be modified for a continuous $\Gamma$ and combined with a covering argument to obtain the bound accordingly.
@@ 5 Numerical Experiments and Discussions
We make the following observations on the numerical experiment. The transformer achieves a better performance than the structured algorithms for each task. This result is impressive to us in that the transformer is not replicating/imitating the performance of these benchmark algorithms but it discovers a new and more effective algorithm. This in fact shows the effectiveness of using the optimal actions $a_{t}^{*}$ as the target variable for training the transformer. Moreover, we attribute the advantage of the transformer against the benchmark algorithms to two aspects: prior knowledge from pre-training data and more greedy exploitation. First, the transformer well utilizes the pre-training data and can be viewed as tailored for the pre-training distribution. Comparatively, all the benchmark algorithms are cold-start. Second, the benchmark algorithms are all designed in ways to cater for asymptotic optimality and thus may sacrifice short-term rewards. Comparatively, the transformer behaves more greedy; just like $\texttt{Alg}^{*}$ is a decision function that does not explicitly encourage exploration. And such greedy behavior brings better reward when $T$ is small such as $T\leq 100.$
We provide more numerical experiments to Section C. In Section C.1, we present additional experiments on training dynamics, including an ablation study on $f$ and $\kappa$ as well as further results extended from Figure 2 (a). Section C.2 provides more results regarding the alignment of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$. Section C.3 compares the performance of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$, showing that $\texttt{TF}_{\hat{\theta}}$ can outperform $\texttt{Alg}^{*}$ in certain scenarios. In Section C.4, we demonstrate that $\texttt{TF}_{\hat{\theta}}$ can be a solution to model misspecifications. We also evaluate the performance of $\texttt{TF}_{\hat{\theta}}$ under environments with varying levels of problem complexity (Section C.5), longer testing horizons (Section C.6), and the case of model misspecification when the test environment is sampled from a distribution different from pre-training (i.e., its OOD performance) (Section C.7). Finally, we conduct an ablation study on the model architecture, model size, and task complexity in Section C.8 and Section C.9.
6 Conclusion
In this paper, we study the pre-trained transformer for sequential decision-making problems. Such problems enable using the optimal action as the target variable for training the transformer. We mathematically formulate the setup of the training pipeline and propose a way to utilize simulation environments to generate pre-training data. We investigate the training and generation aspect of the transformer and identify an OOD issue that is largely ignored by the existing literature. Theoretically, we interpret how the transformer works as an online algorithm through the lens of the Bayes-optimal decision function. The numerical experiments are encouraging in that the trained transformer learns a decision function significantly better than all the benchmark algorithms.
Appendix A Related Works
A.1 Literature Review
@@
A remarkable characteristic of pre-trained transformers is their ability to perform in-context learning (ICL). Once pre-trained on a vast corpus, transformers can solve new tasks with just a few examples, without updating their parameters. ICL has captured the attention of the theoretical machine learning community, leading to considerable efforts into understanding ICL from different theoretical perspectives [Xie et al., 2021, Akyürek et al., 2022, Von Oswald et al., 2023, Lin et al., 2023]. A stream of literature that is related to our work explains the transformer’s behavior as Bayesian inference. See Xie et al. [2021], Zhang et al. [2023b] for the Bayesian argument in regression context, Müller et al. [2022] for approximating a large set of posteriors, and Lee et al. [2024] for the Bayesian analysis in decision-making. Another stream of literature aims to develop generalization bound or other types of convergence results. See Jeon et al. [2024], Bai et al. [2024] for such results in the regression context and Lin et al. [2023] in the decision-making context.
Pre-trained transformer for RL.
The earliest effort of applying transformers to reinforcement learning lies in the area of offline RL [Janner et al., 2021, Chen et al., 2021]. By autoregressively maximizing the likelihood of trajectories in the offline dataset, this paradigm essentially converts offline RL to a supervised learning problem. When evaluating the policy, the actions are sampled according to the likelihood of PT conditioned on the so called “return-to-go”, which is commonly chosen to be the cumulative reward of “good” trajectories in the dataset. By doing so, PT retrieves information from trajectories in training data with similar return-to-go values and performs imitation learning to high-reward offline trajectories. Based on this observation, several works further examine the validity of return-to-go conditioning [Emmons et al., 2021, Ajay et al., 2022, Brandfonbrener et al., 2022] and propose alternative methods that improve this approach or extend to other settings [Laskin et al., 2022, Badrinath et al., 2024, Zheng et al., 2022, Wu et al., 2024]. For example, one of the drawbacks of the return-to-go conditioning is that it often fails in trajectory-stitching when offline data comes from sub-optimal policies. One recent work [Park et al., 2024] finds that pre-trained LLMs, although trained by text data, can behave like a follow-the-perturbed-leader (FTPL) algorithm in bandits and games. It further introduces a regret-based training loss, through which transformers can be trained in an unsupervised way.
@@
A.2 $\texttt{TF}_{\hat{\theta}}$ v.s. Online Decision Transformer
In this section, we compare $\texttt{TF}_{\hat{\theta}}$ with online decision transformer (ODT) [Zheng et al., 2022]. As shown in Figure 5, although both approaches involve a pre-training phase and an (online) interaction phase, they are indeed very different and designed to address different problems. Specifically:
  • ODT interacts with a specific testing environment to fine-tune the policy, whereas $\texttt{TF}_{\hat{\theta}}$ (with Algorithm 1) interacts with different environments and does the learning in the pre-training stage. As a result, ODT is tailored to optimize performance in a single test environment, while $\texttt{TF}_{\hat{\theta}}$ learns various policies across different environments during pre-training, enabling it to perform well in online evaluations across different environments.
  • ODT has been shown to underperform in stochastic environments. Although some RL environments—such as Atari and Gym—are relatively less stochastic, sequential decision-making often involves more stochasticity. This is why ODT has rarely been applied to sequential decision-making problems like bandit problems. More specifically, ODT, along with its precursor Decision Transformer (DT) [Chen et al., 2021], belongs to the class of “return-conditioned supervised learning” methods, which are not suitable even for offline reinforcement learning in stochastic environments. In such cases, the transformers (ODT or DT) is prone to learning from a “survival bias”, as demonstrated in Brandfonbrener et al. [2022], leading to suboptimal performance.
(a) Online Decision Transformer (ODT)
(b) $\texttt{TF}_{\hat{\theta}}$ with Algorithm 1 (ours)
Figure 5: Comparison between training frameworks of Online Decision Transformer (ODT) [Zheng et al., 2022] and $\texttt{TF}_{\hat{\theta}}$. The online interactions of ODT is in the (single) test environment to explore a (single) good policy in the same environment, while our proposed training framework (Algorithm 1) has online interactions in multiple environments to mitigate the OOD issue in the pre-training data (as discussed in Section 3), and $\texttt{TF}_{\hat{\theta}}$ can handle different test environments.
We further conduct an experiment comparing our framework with ODT as shown in Figure 6 to highlight their differences. Specifically, we use a noiseless linear bandits task (i.e., no reward noise) with dimension $d=2$ and a horizon of $T=20$, where the test environments are entirely unseen during the training phase, including the online fine-tuning phase for ODT. It is important to emphasize that this problem setting differs from the one for which ODT is designed (as shown in Figure 5(a), where the test environment is identical to the environment used during online fine-tuning), and aligns with the setting we study in this work. We set fair configurations for training ODT and $\texttt{TF}_{\hat{\theta}}$ (which will be specified later). As expected, Figure 6 demonstrates that ODT fails in this setting, performing no better than a random policy that uniformly samples actions.
Figure 6: We test the ODT and our method in a same linear bandits task, where test environments are unseen before the evaluation phase. The ODT fails in the test environments, which are different from the online fine-tuning phase’s.
Setup: for both ODT and $\texttt{TF}_{\hat{\theta}}$, we set the transformer architecture with 4 layers, 4 heads, 512 embedding dimensions, and a 20-length context window, which are the same as in Zheng et al. [2022]. For the offline dataset, we sample 100,000 environments by the same method in our paper, and then use Thompson sampling to generate trajectories for each environment. And for the online interactions, the environments are sampled from the same distribution as the offline data. The testing environments are still sampled from the same distribution. We set the offline training phase (pre-training stage for ODT and stage 1 for $\texttt{TF}_{\hat{\theta}}$) with 20 iterations (each iteration has 500 updates/gradient descents on parameters), and set the online interaction phase (online fine-tuning for ODT and stage 2 for $\texttt{TF}_{\hat{\theta}}$) with 30 iterations (each iteration has 50 updates/gradient descents on parameters). The other training parameters (e.g., learning rate) are set the same as in Zheng et al. [2022] except the return-to-go (RTG): we set the eval RTG as 55, which is the maximum RTG observed in the offline dataset, and online RTG as 110, which doubles the eval RTG as used in Zheng et al. [2022].
@@ Appendix B Problem Examples for the General Setup in Section 2
For the general problem setup in Section 2, the following proposition relates the pre-training and the regret through the lens of loss function.
@@
$\mathrm{Regret}(f;\gamma)\leq C\cdot L\left(f;\gamma\right)$
holds for any $f$ and $\gamma$. We have
@@
Proposition B.2 gives a first validity of the supervised pre-training approach. The surrogate property is a property of the loss function and it is not affected by the distribution or the decision function $\texttt{TF}_{\theta}.$ Specifically, it upper bounds the regret with the action prediction loss. It is also natural in that a closer prediction of the optimal action gives a smaller regret. One implication is that the result guides the choice of the loss function for different underlying problems.
@@
To prove Proposition B.2, we assume the following mild conditions. First, for all these problems, the noise random variable $\epsilon_{t}$ has mean 0, is i.i.d. across time, and is independent of the action $a_{t}$. Second, for the linear bandits problem, the $L_{\infty}$ norm of the unknown parameter $w$ is bounded by $D$. Third, for the dynamic pricing problem, we assume the demand function is differentiable with respect to the action coordinate, and both the absolute value of the first- and second-order derivatives is bounded by $D$. Further details regarding the problem formulations can be found in Appendix E.2.
@@
Since both the regret $\text{Regret}(f;\gamma)$ and the prediction error $L(f;\gamma)$ are the sums of single-step regret and prediction error over the horizon, it is sufficient to prove the single-step surrogate property of the loss function. Therefore, in the following proof, we will focus on proving the single-step surrogate property and omit the subscript $t$ for simplicity.
  • For the multi-arm bandits problem, without loss of generality, we assume $\mathcal{A}=\{a_{1},\ldots,a_{J}\}$ and the optimal arm is $a_{1}$. For $1\leq i\leq J$, let $\mu_{j}$ be the mean of the reward distribution for arm $j$, and define $\Delta_{j}=\mu_{1}-\mu_{j}$ as the reward gap for arm $j$. Define $\Delta_{\max}=\max\left\{\Delta_{1},\cdots,\Delta_{J}\right\}$ as the maximum action gap. Under cross-entropy loss, we denote the output distribution of decision function $f$ is $(p_{1},\ldots,p_{J})$. Then the single-step regret can be bounded by:
    $\mathbb{E}\left[r\left(X,a^{*}\right)-r\left(X,f(H)\right)|H\right]=\mu_{1}-\sum_{j=1}^{J}p_{j}\mu_{j}=\sum_{j=2}^{J}p_{j}\Delta_{j}\leq\Delta_{\max}(1-p_{1})\leq-\Delta_{\max}\cdot\log(p_{1}),$
    where the last step is by the inequality $1-x\leq-\log(x)$ for $0<x\leq 1$. Then by the definition of cross-entropy loss, we finish the proof.
  • For the dynamic pricing problem, we have
    $r(X,a^{*})-r(X,f(H))=\nabla_{a}r(X,a^{*})(a^{*}-f(H))+\nabla_{a}^{2}r(X,\tilde{a})|f(H)-a^{*}|^{2}\leq D|f(H)-a^{*}|^{2}$
    where $\tilde{a}$ is in the line of $f(H)$ and $a^{*}$. Here, the first step is by the Taylor expansion, and the second step is by the first order condition of $a^{*}$ and the assumption. Then by noting $|f(H)-a^{*}|^{2}$ is the square loss, we finish the proof.
  • For the linear bandits problem, we have
    $r(X,a^{*})-r(X,f(H))=w^{\top}a^{*}-w^{\top}f(H)\leq||w||_{\infty}||f(H)-a^{*}||_{1}\leq D||f(H)-a^{*}||_{1},$
    where the second step is by the Holder’s inequality and the last step is by the assumption. Then by noting $||f(H)-a^{*}||_{1}$ is the absolute loss, we finish the proof.
  • For the newsvendor problem, we have
    $r(X,a^{*})-r(X,f(H))$ $\leq\max\left\{h,l\right\}\mathbb{E}_{X}\left[|a^{*}-d(X)-f(H)+d(X)|\right]$
    $=\max\left\{h,l\right\}|a^{*}-f(H)|$
    where the first line is because the (random) reward function is $\max\left\{h,l\right\}$-Lipschitz in $a-d(X)$. Then by noting $|f(H)-a^{*}|$ is the absolute loss, we finish the proof.